OpenStudy (anonymous):

can me explain the procedure to resolve? 1/a + 1/2 = 2/a

jimthompson5910 (jim_thompson5910):

You can multiply both sides by the LCD 2a to get... \[\Large \frac{1}{a} + \frac{1}{2} = \frac{2}{a}\] \[\Large 2a\left(\frac{1}{a} + \frac{1}{2}\right) = 2a\left(\frac{2}{a}\right)\] \[\Large 2a\left(\frac{1}{a}\right) + 2a\left(\frac{1}{2}\right) = 2a\left(\frac{2}{a}\right)\] \[\Large \frac{2a}{a} + \frac{2a}{2} = \frac{4a}{a}\] \[\Large \frac{2\cancel{a}}{\cancel{a}} + \frac{\cancel{2}a}{\cancel{2}} = \frac{4\cancel{a}}{\cancel{a}}\] \[\Large 2 + a = 4\] Notice how those pesky fractions have been eliminated after multiplying every term by the LCD.

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

What solution do you get?

OpenStudy (anonymous):

A= 2 ?

jimthompson5910 (jim_thompson5910):

Check: \[\Large \frac{1}{a} + \frac{1}{2} = \frac{2}{a}\] \[\Large \frac{1}{2} + \frac{1}{2} = \frac{2}{2}\] \[\Large \frac{1+1}{2} = \frac{2}{2}\] \[\Large \frac{2}{2} = \frac{2}{2} \ \ {\color{green}{\checkmark}}\] so a = 2 is definitely the answer

jimthompson5910 (jim_thompson5910):

it's always a good idea to check the answer back into the ORIGINAL equation because sometimes you'll get potential solutions that don't work at all (they are called extraneous solutions)

OpenStudy (anonymous):

ok thanks you

jimthompson5910 (jim_thompson5910):

np

OpenStudy (loser66):

 just contribute another way to solve